A grocer wants to make a mixture of three dried fruits. He decides that the ratio of pounds of banana chips to apricots to dates should be 3:1:1. Banana chips cost $1.17/lb, apricots cost $3.00/lb, and dates cost $2.30/lb. What is the cost per pound of the mixture?
I have tried to organize the info into a table but i still dont seem to get it and I need help formulating an equation.
Please help, thanks.

Answer by stanbon(75887)

A grocer wants to make a mixture of three dried fruits. He decides that the ratio of pounds of banana chips to apricots to dates should be 3:1:1. Banana chips cost $1.17/lb, apricots cost $3.00/lb, and dates cost $2.30/lb. What is the cost per pound of the mixture?
-------
3:1:1 is the same as 3x:x:x
-----
1.17(3x) + 3x + 2.3x = 8.81
-----
Average price = 8.81/(3+1+1) = $1.76/lb
===================
Cheers,
Stan H.
===================